How long would it take to travel 80 miles at 80 miles per hour?
Found 2 solutions by Alan3354, mathhelp@:
Answer by Alan3354(30966)

How long would it take to travel 80 miles at 80 miles per hour?
-------------
d = r*t
t = d/r

Answer by mathhelp@(10)

distance = time * speed
so your distance is 80 miles
speed is 80 miles per hour
so your question is asking how much time (x)
80 = x * 80
x is 1 hour
:P